OpenStudy (anonymous):

3(x-3)=6x-18 ------ 2

OpenStudy (anonymous):

1st distribute

OpenStudy (anonymous):

and Welcome to OpenStudy :)

OpenStudy (compassionate):

\[3(x - 3) = \frac{ 6x - 18 }{ 2 } \rightarrow 3x - 9 = 3x - 9\] Basically what I di was distribute the 3. On the right side, I divided the entire equation by the Highest Common Denominator, which is 2. Giving you the equation: 3x - 9 = 3x - 9 Now, just solve like a normal linear equation.

OpenStudy (anonymous):

so I solved it but got no solution. not sure if that's rights??

OpenStudy (compassionate):

Correct. There is no solution. 3x - 9 = 3x - 9 3x - 9 + 9 = 3x - 9 + 9 3x = 3x x = x 0

OpenStudy (anonymous):

awesome thanks

OpenStudy (compassionate):

No problem. Come back to OpenStudy. We're willing to help you! It is late at night, so not many people are online. If you're having trouble, feel free to message or tag people in post (Tag by adding an @ symbol + their name)

OpenStudy (anonymous):

yeah I've got lots to do. I'm hoping you can probably help me I would really appreciate it

OpenStudy (compassionate):

Feel free to tag me in any problems you may have. Remember, don't ask more than one question in a post. If you have a new question, open a new topic.

OpenStudy (compassionate):

I'll be AFK for a few minutes. Going to grab a bite to eat.